OpenStudy (anonymous):

The sum of the infinite series 1/5 - 2/5^2 + 2^2/5^3 - 2^3/5^4 + ...... is

hartnn (hartnn):

the common ratio is -2/5, isn't it ?

OpenStudy (anonymous):

Be smart

OpenStudy (anonymous):

it is a geometric series, use the well known formula for it's sum

hartnn (hartnn):

First term = 1/5 so sum to infinity = (1/5)/(1-(-2/5))=(1/5)/(7/5)=1/7.

OpenStudy (anonymous):

What is the common ratio if it is GP?

OpenStudy (anonymous):

@hartnn said it already

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

the common ratio is -2/5, isn't it ? that what first answer says.

hartnn (hartnn):

Actually it is -2/5...why it is not ?? @Zekarias

OpenStudy (anonymous):

\[(\frac{-2}{5})^{0}=1\]

OpenStudy (anonymous):

Ok if the common ratio is that and the first term is 1/5 write me 1/5 in terms of -2/5

OpenStudy (anonymous):

read what i just wrote

OpenStudy (anonymous):

But the first term is 1/5 not 1

OpenStudy (anonymous):

1/5*(-2/5)^0=1/5

hartnn (hartnn):

WHY do u want to do that ? there is no reason to express 1/5 in terms of -2/5 to get the sum....the sum to infinity is still 1/7.....

OpenStudy (anonymous):

=>1/5=2^0/5^1

OpenStudy (anonymous):

correct all of you. Thanks

OpenStudy (anonymous):

the general term of this series is: \[a*q ^{n}\] where a=1/5 , q=-2/5 and n goes fomr n=0 to infinity